Difference between mirror reflected light rays and rays of a screen

Grade:9

1 Answers

Saurabh Koranglekar
3 years ago
Dear student

Distance to the image is the big difference.

The reflected image in a plane mirror is composed of rays that appear to come from a pointbehindthe mirror (by the same distance that the image that is being reflected is in front of the mirror:d0=did0=di). For non-plane mirrors you'll have to work a little harder, but locally spherical bits can be treated with the usual1/f=1/di+1/do.1/f=1/di+1/do.The light rays emanating from a screen come from the surface of the screen.

Regards